if the lenghts of two segments of a foal chord intercepted by its axis are 3 and 5 then the lengt of the latus rectum would be?

Grade:12

1 Answers

Arun
4 years ago
Dear student
 
Length of latus rectum = 2 bc/(b+c)
when b and c are lengths of focal chord intercepted by axis
 
hence length of latus rectum = 2 * 3*5/(3+5) = 30/8